# SAM2-Adapter: Evaluating & Adapting Segment Anything 2 in Downstream Tasks: Camouflage, Shadow, Medical Image Segmentation, and More

### v1

@@ 2 Related Work
Semantic Segmentation. In recent years, semantic segmentation has made significant progress, primarily due to the remarkable advancements in deep-learning-based methods such as fully convolutional networks (FCN) [49], encoder-decoder structures [62, 75, 21, 1, 42, 11], dilated convolutions [5, 6, 7, 8, 48, 76], pyramid structures [89, 7, 79, 8, 91, 25], attention modules [78, 24, 96, 94, 92], and transformers [82, 72, 67, 14, 93]. Recent advancements have improved SAM’s performance, such as [35], which introduces a High-Quality output token and trains the model on fine-grained masks. Other efforts have focused on enhancing SAM’s efficiency for broader real-world and mobile use, exemplified by [74, 77, 80]. The widespread success of SAM has led to its adoption in various fields, including medical imaging [51, 17, 53, 71], remote sensing [4, 61], motion segmentation [73], and camouflaged object detection [68]. Notably, our previous work SAM-Adapter [10, 9] tested camouflaged object detection, polyp segmentation, and shadow segmentation, and provide the first adapter-based method to integrate the SAM’s exceptional capability to these downstream tasks.
Adapters. The concept of Adapters was first introduced in the NLP community [29] as a tool to fine-tune a large pre-trained model for each downstream task with a compact and scalable model. In [66], multi-task learning was explored with a single BERT model shared among a few task-specific parameters. In the computer vision community, [43] suggested fine-tuning the ViT [18] for object detection with minimal modifications. Recently, ViT-Adapter [13] leveraged Adapters to enable a plain ViT to perform various downstream tasks. [47] introduce an Explicit Visual Prompting (EVP) technique that can incorporate explicit visual cues to the Adapter. However, no prior work has tried to apply Adapters to leverage pretrained image segmentation model SAM trained at large image corpus. Here, we mitigate the research gap.

### v2

@@ 2 Related Work
Semantic Segmentation. In recent years, semantic segmentation has made significant progress, primarily due to the remarkable advancements in deep-learning-based methods such as fully convolutional networks (FCN) [56], encoder-decoder structures [70, 19, 1, 41, 9, 37, 35], dilated convolutions [5, 6, 55, 86, 28], pyramid structures [98, 5, 88, 6, 100, 23, 34, 81], attention modules [105, 103, 101, 40, 36], and transformers [91, 83, 75, 12, 102]. Recent advancements have improved SAM’s performance, such as [43], which introduces a High-Quality output token and trains the model on fine-grained masks. Other efforts have focused on enhancing SAM’s efficiency for broader real-world and mobile use, exemplified by [85, 87, 89]. The widespread success of SAM has led to its adoption in various fields, including medical imaging [58, 15, 61, 82, 59], remote sensing [4, 69, 38], motion segmentation [84, 79, 78, 21], and camouflaged object detection [76]. Notably, our previous work SAM-Adapter [8, 7] tested camouflaged object detection, polyp segmentation, and shadow segmentation, and provide the first adapter-based method to integrate the SAM’s exceptional capability to these downstream tasks.
Adapters. The concept of Adapters was first introduced in the NLP community [27] as a tool to fine-tune a large pre-trained model for each downstream task with a compact and scalable model. In [74], multi-task learning was explored with a single BERT model shared among a few task-specific parameters. In the computer vision community, [50, 39, 33] suggested fine-tuning the ViT [16] for object detection with minimal modifications. Recently, ViT-Adapter [11] leveraged Adapters to enable a plain ViT to perform various downstream tasks. [54] introduce an Explicit Visual Prompting (EVP) technique that can incorporate explicit visual cues to the Adapter. However, no prior work has tried to apply Adapters to leverage pretrained image segmentation model SAM trained at large image corpus. Here, we mitigate the research gap.